OpenStudy (anonymous):

derivative of f(x)=csc(pi*x^2). Can't you just start w/ f'(x)= -csc(pi*x^2)*cot(pi*x^2) ?

OpenStudy (abb0t):

\[\sf \csc(\pi x^2)~??\]

OpenStudy (anonymous):

yep

OpenStudy (abb0t):

Are you familiar with chain rule? derivative of the outside times the dervative of the inside. In otherwords, you have: \(\sf \color{red}{\frac{d}{dx}outside(inside) \times \frac{d}{dx}(inside)}\)

OpenStudy (abb0t):

Remembr that \(\pi\) is a constant.

OpenStudy (anonymous):

I mean, isn't the following true? The derivative of: \[\csc(x) = -\csc(x)\cot(x)\]

OpenStudy (anonymous):

I'm familiar w/ the chain rule but it's getting a little fuzzy w/ the trig functions

OpenStudy (abb0t):

Yes, thats correct. So you would have\(\sf \color{blue}{-cot(\pi x^2)(csc(\pi x^2)}\)\(\sf \color{red}{\times \frac{d}{dx}(\pi x^2)}\) = ____________________

OpenStudy (anonymous):

ok, so I can't just take the derivative like I did 3 lines up then? It's still a chain rule type relationship w/ these?

OpenStudy (abb0t):

I don't know what you mean. Unless you're asked to find \(\sf \color{purple}{f''(x)}\), what I showed up there, is all you need to do. And of course, don't forget to take the deriviative of the inside function, as I did not finish the problem. I left it for you to finish.

OpenStudy (anonymous):

ok, so take a more generic approach. let \[f(x)=\csc(x)\] \[f \prime(x)=-\csc(x)\cot(x)\] or is: \[f \prime(x)=-\csc(x)\cot(x)*(x)'\] Sorry, I'm just learning the syntax.

OpenStudy (abb0t):

Yes.

OpenStudy (anonymous):

Which one? The first or the second solution?

OpenStudy (abb0t):

The second one.

OpenStudy (anonymous):

ok, somehow I missed that in my notes. I thought it was just the first.

OpenStudy (abb0t):

According to the chain rule, the second one is correct, but you don't need to show that since you know that \(\sf \frac{d}{dx}~x = 1\)

OpenStudy (anonymous):

oh, I see now... in class they were just using sin(x) (etc.) in the examples. They didn't do things like sin(pi*x) and elaborate on that.

OpenStudy (anonymous):

That makes sense now

OpenStudy (anonymous):

Thanks so much. You rock and this site is pretty baller!